# Engineering subgap states in superconductors by altermagnetism

### v3

@@ Abstract
We investigate the realization and control of subgap states by tailored altermagnetic fields on unconventional superconductors. When the symmetries of altermagnetism and unconventional superconductivity align, we demonstrate the emergence of bulk zero-energy flat bands, giving rise to a zero-bias conductance peak. The symmetry and strength of $d$- and $g$-wave altermagnets strongly affect the surface Andreev states from $d$-wave and chiral $d$- and $p$-wave superconductors. As a result, distinct types of subgap states are realized, including curved and flat bands, that can be detected by tunneling spectroscopy. Furthermore, we find that the altermagnetism-induced subgap states give rise to a large spin conductance at zero net magnetization which helps identify the strength of the underlying altermagnetism and superconductivity. Our results offer a solid route for designing and manipulating subgap states in superconducting systems, which can be useful for functionalizing superconducting spintronic devices.
@@
Figure 1: An altermagnet (AM, orange) coupled to a superconductor (SC, green) is used to engineer subgap states. A normal metal probe (N, gray) coupled to the AM-SC system at $x=0$ is then used to probe the induced subgap states via differential conductance $dI/dV$. (b) Table showing the types of emergent subgap states with respect to $k_{y}$ induced by altermagnetism in distinct unconventional SCs.
@@
In this work, we demonstrate the emergence and control of highly tunable subgap states by using altermagnetism in unconventional SCs. To show this, we focus on a hybrid system composed of a two-dimensional SC with altermagnetism due to its proximity to an AM [Fig. 1(a)], where a normal probe enables the detection of bulk and surface subgap states. We consider various types of unconventional SCs combined with $d$- and $g$-wave AMs, which allows us to reveal various types of subgap states, including bulk zero energy flat bands and SABSs and listed in Fig. 1(b). We further demonstrate that these subgap states give rise to strong charge and spin conductance signatures, enabling their detection and also the identification of the type of altermagnetism supporting them. Our findings demonstrate that altermagnetism can be effectively harnessed to generate tunable subgap states and spin currents in superconducting systems.
Model.— We aim to harness altermagnetism to realize and detect subgap states in unconventional SCs. We thus consider a junction along $x$, where an unconventional SC coupled to a two-dimensional AM is connected to a normal metal lead N at $x=0$, see Fig. 1(a). The AM induces an altermagnetic order parameter into the SC via proximity effect and the lead can probe surface and bulk properties. The altermagnetic superconducting junction is modeled by $\mathcal{H}=(1/2)\sum\nolimits_{\bm{k}}\hat{c}_{\bm{k}}^{{\dagger}}\mathcal{H}_{\bm{k}}\hat{c}_{\bm{k}}$, where the Nambu spinor is given by $\hat{c}_{\bm{k}}=(c_{\bm{k}\uparrow},c_{\bm{k}\downarrow},c_{-\bm{k}\uparrow}^{{\dagger}},c_{-\bm{k}\downarrow}^{{\dagger}})^{T}$, $\bm{k}=(k_{x},k_{y})$, and
@@
[…]
@@
where $\nu=\pm$ indicating two possible anisotropic energy shifts ($\pm h_{0}m_{\bm{k}}$) of subbands. From Eq. (2), it is easy to see that, for $\nu h_{0}m_{\bm{k}}=|\Delta_{0}s_{\bm{k}}|$, we get $E^{+}_{\bm{k}\nu}\neq 0$ and $E^{-}_{\bm{k}\nu}=0$, the latter signaling a zero energy bulk flat band akin to BFSs [23]. This is an example of a bulk subgap state induced by altermagnetism, and its detection can be carried out via conductance in the setup of Fig. 1(a). Notably, this setup also allows us to search for altermagnetism-induced subgap states at the interface surface at $x=0$, which can be directly accessed via conductance as we explain below.
To obtain the conductance, we employ the scattering approach, which is based on the wavefunctions of injected electrons with spin $\sigma=\uparrow,\downarrow$. Due to translational invariance along $y$, the wavefunctions are $\Psi_{\sigma}(x,k_{y})=\psi_{\sigma}(x)e^{iyk_{y}}$. In N, $\psi_{\sigma}(x)$ is given by
$\psi_{\sigma}(x<0)=[\hat{A}_{\sigma}+a_{\sigma}\hat{B}_{\sigma}]e^{ixk_{x}}+b_{\sigma}\hat{A}_{\sigma}e^{-ixk_{x}}\,$, (3)
where $\hat{A}_{\uparrow}=\left(1,0,0,0\right)^{T}$, $\hat{A}_{\downarrow}=\left(0,1,0,0\right)^{T}$, $\hat{B}_{\downarrow}=\left(0,0,1,0\right)^{T}$ and $\hat{B}_{\uparrow}=\left(0,0,0,1\right)^{T}$, $k_{x}=k_{F}\cos\theta$, and $k_{y}=k_{F}\sin\theta$. The wave functions in the SC can be written in a similar way. Then, the coefficients $a_{\sigma}$ and $b_{\sigma}$ are obtained from the boundary conditions $\psi(0^{+})=\psi(0^{-})$, $\partial_{x}\psi(0^{+})-\partial_{x}\psi(0^{-})=(2mU_{0}/\hbar^{2})\psi(0)$. The angle-resolved charge (spin) conductance is found as [98, 99, 100] $\bar{\sigma}_{\sigma}^{e(s)}=1\pm|a_{\sigma}|^{2}-|b_{\sigma}|^{2}$, which gives
$\bar{\sigma}_{\sigma}^{e(s)}=\bar{\sigma}_{n}\frac{1\pm\bar{\sigma}_{n}|\Gamma_{\sigma}^{+}|^{2}+(\bar{\sigma}_{n}-1)|\Gamma_{\sigma}^{+}\Gamma_{\sigma}^{-}|^{2}}{|\Lambda-\eta(1-\bar{\sigma}_{n})\Gamma_{\sigma}^{+}\Gamma_{\sigma}^{-}|^{2}}\,$, (4)
@@
where $E_{\uparrow,\bm{k}_{\pm}}=E-\mathcal{M}_{\bm{k}_{\pm}}$ and $E_{\downarrow,\bm{k}_{\pm}}=E+\mathcal{M}_{\bm{k}_{\pm}}$, with $\bm{k}_{+}$, $\bm{k}_{-}$ are given by $\bm{k}_{\pm}=k_{F}\left(\pm\cos\theta,\sin\theta\right)$. Moreover, $\Lambda=1$ for $d_{x^{2}-y^{2}}$- and $d_{xy}$-wave SCs; $\Lambda=e^{2i\theta}$ for chiral $p$-wave SCs; $\Lambda=e^{4i\theta}$ for chiral $d$-wave SCs. Also, $\eta=1$ for spin-singlet SCs (cases i,ii,iv), while $\eta=-1$ for spin-triplet SCs (case iii). Hence, the charge (spin) conductance in the superconducting state is calculated as
$\sigma_{\rm SC}^{e(s)}=\frac{e^{2}}{h}\int_{-\frac{\pi}{2}}^{\frac{\pi}{2}}\bar{\sigma}^{e(s)}\cos\theta d\theta.$ (6)
where $\bar{\sigma}^{e}=\bar{\sigma}_{\uparrow}^{e}+\bar{\sigma}_{\downarrow}^{e}$ and $\bar{\sigma}^{s}=\bar{\sigma}_{\uparrow}^{s}-\bar{\sigma}_{\downarrow}^{s}$ are the angle-resolved charge and spin conductances, respectively. $\sigma_{\rm SC}^{s}$ measures the ability to produce a spin current and plays a key role in superconducting spintronics. Later we normalize the conductances by $\sigma_{N}=(2e^{2}/h)\int\bar{\sigma}_{n}\cos\theta d\theta$.
@@
Bulk flat bands by $d_{x^{2}-y^{2}}$-wave altermagnetism in $d_{x^{2}-y^{2}}$-wave SCs.— We begin by analyzing the junction formed by $d_{x^{2}-y^{2}}$-wave SC with $d$-wave altermagnetism. Before delving into the conductance, we remind our discussion below Eq. (2), where we show that at $\nu h_{0}m_{\bm{k}}=|\Delta_{0}s_{\bm{k}}|$, one of the bands reaches zero energy, namely, $E^{-}_{\bm{k}\nu}=0$, while $E^{+}_{\bm{k}\nu}\neq 0$. Thus, when the AM and SC have the same symmetry ($m_{\bm{k}}=s_{\bm{k}}$) a zero-energy flat band emerges as a bulk subgap state entirely induced by altermagnetism and akin to BFSs. As a result, the bulk density of states at zero energy becomes singular, similar to the divergent density of states at $E=\Delta_{0}$ for a conventional $s$-wave SC. This effect, and hence the flat band, can be detected via conductance, as we demonstrate in Fig. 2, where we plot the charge (spin) conductance for both $d_{x^{2}-y^{2}}$- and $d_{xy}$-wave AMs with $d_{x^{2}-y^{2}}$-wave superconductivity. The first feature to notice is that, at $h_{0}\neq\Delta_{0}$, the angle-resolved charge conductance $\bar{\sigma}^{e}(eV)$ is greatly enhanced at $eV=0$ as a result of the BFS ($E^{-}_{\bm{k}\nu}=0$) discussed above [Fig. 2(a)]. This gives rise to a zero-bias conductance peak (ZBCP) in the total charge conductance $\sigma_{\rm SC}^{e}(eV)$ at $h_{0}\neq\Delta_{0}$, which already appears at $Z=0$ but is enhanced as $Z$ enhances [Fig. 2(b)]. Thus, the ZBCP directly results is a bulk effect due to the BFS and it is not related to SABSs. By moving away from $eV=0$ and for $h_{0}\neq\Delta_{0}$, $\sigma_{\rm SC}^{e}(eV)$ develops a V-shape profile with spikes outside the gap [Fig. 2(c)], an effect that is similar to ferromagnetic SCs [39]. On the other hand, when $d_{xy}$-wave AM and $d_{x^{2}-y^{2}}$-wave SC coexist, no zero energy bulk flat band is found although the charge conductance still acquires finite subgap values following a U-shape profile for $h_{0}\neq 0$ but without any ZBCP, see Fig. 2(e,f).
To further unveil the consequences of the ZBCP and BFS found here, in Fig. 2(d) we show the spin conductance $\sigma_{\rm SC}^{s}(eV)$ found from Eqs. (6) as a function of $eV$ and distinct $h_{0}$ for a $d_{x^{2}-y^{2}}$-wave AM and a $d_{x^{2}-y^{2}}$-wave SC. For $h_{0}\neq 0$, a $\sigma_{\rm SC}^{s}(eV)\neq 0$ appears reflecting the role of $d_{x^{2}-y^{2}}$-wave altermagnetism, where $\Gamma_{\uparrow}^{\pm}\neq\Gamma_{\downarrow}^{\pm}$ [Eq. (5)]. We further see that the spin conductance bears multiple sign changes as $eV$ varies, and, interestingly, it exhibits an abrupt sign change for $h_{0}=\Delta_{0}$ at $eV=0$. Thus, the abrupt sign change in $\sigma_{\rm SC}^{s}(eV=0)$ is a direct consequence of the zero-energy bulk flat band when the symmetry of the AM and SC are the same. For distinct symmetries, as for a $d_{xy}$-wave AM and a $d_{x^{2}-y^{2}}$-wave SC in the inset of Fig. 2(f), the spin conductance vanishes due to absence of spin splitting for $d_{xy}$-wave AMs leading to $\Gamma_{\uparrow}^{\pm}=\Gamma_{\downarrow}^{\mp}$. Hence, having AMs and SCs with the same symmetry enables the realization of bulk zero-energy flat bands that can be detected via charge conductance and can be used to efficiently manipulate spin currents in an electric way without a net magnetization [102, 103, 104].
Figure 2: (a-d) Conductance for a junction having a $d_{x^{2}-y^{2}}$-SC with a $d_{x^{2}-y^{2}}$-AM (a-d), including the angle-resolved charge conductance at $Z=2$, $h_{0}=\Delta_{0}$ (a), the total zero-bias charge conductance as a function of $Z$ and $h_{0}$ (b), the total charge conductance as a function of eVeV for distinct $h_{0}$ at $Z=2$ (c), and the total spin conductance (d). (e,f) The same as in (a,c) for a $d_{xy}$-AM, with the inset in (f) its total spin conductance.
@@
Surface curved and flat bands by $d_{xy}$-wave altermagnetism in $d_{xy}$-wave SCs.— It is well-known that $d_{xy}$-wave pair potentials lead to zero-energy surface flat states without AMs [12]. Here we are interested in analyzing how altermagnetism can control the surface states and how to induce them in regimes not present before. This is demonstrated in Fig. 3, where we plot the angle-resolved and total charge conductances. For a $d_{x^{2}-y^{2}}$-AM and $d_{xy}$-wave SC, the angle-resolved charge conductance $\bar{\sigma}^{e}$ as a function of $\theta$ shows that the surface states split in a curved shape, with the size of the splitting given by the anisotropic AM field, but remain degenerate at $\theta=\pi/4$ [Fig. 3(a)]. Due to this reason, the total charge conductance $\sigma_{\rm SC}^{e}$ gets suppressed as the AM field increases [Fig. 3(b,c)], going from a sharp ZBCP at $h_{0}=0$ into an almost constant subgap conductance for $h_{0}>0$ whose size determines the strength of altermagnetism. Fig. 3(c) further uncovers that this surface effect needs $Z>0$.
Figure 3: (a-c) Conductance for a junction having a $d_{xy}$-SC with a $d_{x^{2}-y^{2}}$-AM, including the angle-resolved charge conductance at $Z=2$, $h_{0}=\Delta_{0}$ (a), the total charge conductance as a function of eVeV for distinct $h_{0}$ at $Z=2$ (b), and the total zero-bias charge conductance as a function of $Z$ and $h_{0}$ (c). (d-f) The same as in (a-c) but for a $d_{xy}$-AM; the inset in (d) shows angle-resolved charge conductance at $h_{0}=1.5\Delta_{0}$.
For a $d_{xy}$-wave AM and $d_{xy}$-wave SC, where the superconducting gap nodes are intact, the surface flat bands inherent to $d_{xy}$-wave SCs remain robust against the AM field when $h_{0}<\Delta_{0}$ but are split for $h_{0}>\Delta_{0}$, see Fig. 3(d) and its inset. This leads to a ZBCP in $\sigma_{\rm SC}^{e}(eV)$ for $h_{0}<\Delta_{0}$, but it reduces when $h_{0}>\Delta_{0}$ into an almost constant value within $|eV|<h_{0}-\Delta_{0}$ [Fig. 3(e,f)]. Further insights are obtained by looking at the zero-bias angle-resolved charge conductance for the $d_{xy}$-AM,
@@
Figure 4: (a,b) Conductance for a junction having a chiral $p$-wave SC with a $d_{x^{2}-y^{2}}$-wave AM, including the angle-resolved charge conductance at $Z=2$, $h_{0}=0.5\Delta_{0}$ (a) and the total charge conductance as a function of eVeV for distinct $h_{0}$ (b). (c,d) The same quantities as in (a,b) but with a $d_{xy}$-AM.
Surface curved bands by $d$-wave altermagnetism in chiral $p$-wave SCs.— Chiral $p$-wave pairing was once believed to describe superconductivity in Sr2RuO4 with chiral edge states [105, 106, 107], and a recent study even supports that Sr2RuO4 supports intrinsic AM order [108]. Motivated by these ideas, here we consider a one-band model with chiral $p$-wave pairing and $d$-wave altermagnetism; in the End Matter, we also address $p$-wave pairing with $g$-wave altermagnetism. For a $d_{x^{2}-y^{2}}$-AM, we calculate the chiral SABSs by imposing $\Gamma_{\sigma}^{+}\Gamma_{\sigma}^{-}=-e^{2i\theta}$ in Eq. (7),
$E_{\text{ABS}\nu}=\Delta_{0}\sin\theta+\nu h_{0}\cos 2\theta\,.$ (9)
with $\nu=\pm$. This reveals that the altermagnet field $h_{0}$ induces a spin splitting in the chiral SABSs $E_{\text{ABS}}^{\pm}$, which appear with a curved dispersion with $\theta$ and also revealed by angle-resolved conductance in Fig. 4(a). The splitting arises due $\Gamma_{\uparrow}^{+}\Gamma_{\uparrow}^{-}\neq\Gamma_{\downarrow}^{+}\Gamma_{\downarrow}^{-}$ in a $d_{x^{2}-y^{2}}$-wave AM. At certain angles found from $\Delta_{0}\sin\theta=\pm h_{0}\cos 2\theta$, however, the SABSs remain degenerate at zero energy [Fig. 4(a)], which affects their dispersion and makes them to acquire a quasi-flat band profile for a range of $\theta$. These quasi-flat bands then induce spikes in the total charge conductance Fig. 4(b) whose separation for $h_{0}>\Delta_{0}/(2\sqrt{2})$ is given by $\Delta_{0}^{2}/\left(4h_{0}\right)+2h_{0}$ from $\partial E_{\text{ABS}\nu}/\partial\theta=0$ and hence determines the strength of altermagnetism.

### v4

# Engineering subgap states in superconductors by the symmetry of altermagnetism

@@ Abstract
Combining superconducting and magnetic materials is a promising path to generate exotic interface subgap states. In this regard, altermagnetism is particularly interesting because it lifts spin degeneracy while providing tailored anisotropy of spin splittings. Here, we investigate the realization and control of subgap states by using the symmetry contrast between altermagnetic fields and unconventional pairings. When the symmetries of altermagnetism and unconventional superconductivity align, we demonstrate the emergence of bulk zero-energy flat bands as the Bogoliubov Fermi surface, giving rise to a zero-bias conductance peak. The symmetry and strength of $d$-wave altermagnets strongly affect the surface Andreev states from $d$-wave and chiral $d$- and $p$-wave superconductors. As a result, distinct types of subgap states are realized, including curved and flat bands, that can be detected by tunneling spectroscopy. Our results offer a solid route for designing and manipulating subgap states in superconducting systems, which can be useful for functionalizing superconducting devices.
@@
Figure 1: (a) An altermagnet (AM, orange) is placed on top of a superconductor (SC, green) to engineer subgap states. A normal metal probe (N, blue) is coupled to SC at $x=0$ near the AM-SC interface, achieved by adjusting the thickness of the insulating substrate (grey) beneath N. (b) Illustrations of the selected AM orderings in momentum space where the red and blue colors indicate up and down spins. (c) Table listing the type of emergent subgap states for several AM-SC combinations. Labels “flat”, “split”, and “curved” indicate when the dispersion of subgap states is respectively zero, spin-split, or drastically varying.
@@
In this work, we demonstrate the emergence and control of highly tunable subgap states by using altermagnetism in unconventional SCs. To show this, we focus on a hybrid system composed of a two-dimensional SC with altermagnetism due to its proximity to an AM [Fig. 1(a)], where a normal probe enables the detection of bulk and surface subgap states. We consider various types of unconventional SCs combined with $d$-wave AMs [Fig. 1(b)], which allows us to reveal different types of subgap states, including bulk zero energy flat bands and SABSs, as listed in Fig. 1(c). Such subgap states are driven by the symmetry contrast between competing superconducting and magnetism orders and give rise to strong charge conductance signatures, enabling their detection and also the identification of the type of altermagnetism supporting them.
Model.— We aim to harness altermagnetism to realize and detect subgap states in unconventional SCs. We thus consider a bilayer AM-SC device where the coupling between AM and SC accounts for a proximity-induced AM field in the SC. Then, a two-dimensional normal metal lead N is connected to the SC at $x=0$ to probe the properties of altermagnetic superconductors, see Fig. 1(a). The N-SC planar junction is modeled by $\mathcal{H}=(1/2)\sum\nolimits_{\bm{k}}\hat{c}_{\bm{k}}^{{\dagger}}\mathcal{H}_{\bm{k}}\hat{c}_{\bm{k}}$, where the Nambu spinor is given by $\hat{c}_{\bm{k}}=(c_{\bm{k}\uparrow},c_{\bm{k}\downarrow},c_{-\bm{k}\uparrow}^{{\dagger}},c_{-\bm{k}\downarrow}^{{\dagger}})^{T}$, $\bm{k}=(k_{x},k_{y})$, and
@@
We consider highly relevant AMs exhibiting $d$-wave parity [Fig. 1(b)] expected in, e.g., RuO2 and KV2Se2 [42, 97], and modeled by specific anisotropic factors $m_{\bm{k}}$. We have $m_{\bm{k}}=(\hat{k}_{x}^{2}-\hat{k}_{y}^{2})\cos 2\alpha+2\hat{k}_{x}\hat{k}_{y}\sin 2\alpha$, with $\alpha=0$ for $d_{x^{2}-y^{2}}$-wave AMs and $\alpha=\pi/4$ for $d_{xy}$-wave AMs. For the unconventional SCs, we consider four types of experimentally relevant pairings [98, 12]. For $d$-wave SCs, we have $\hat{\Delta}_{\bm{k}}=-\Delta_{0}s_{\bm{k}}\hat{s}_{y}\hat{\tau}_{y}$, with $s_{\bm{k}}=(\hat{k}_{x}^{2}-\hat{k}_{y}^{2})\cos 2\chi+2\hat{k}_{x}\hat{k}_{y}\sin 2\chi$. Here, i) $\chi=0$ represents $d_{x^{2}-y^{2}}$-wave pairing and ii) $\chi=\pi/4$ $d_{xy}$-wave. These two pairing states are related by a rotation with respect to the orientation of the interface with the metallic probe. Chiral SCs display iii) chiral $p$-wave pairing $\hat{\Delta}_{\bm{k}}=\Delta_{0}\hat{s}_{x}(\hat{k}_{x}\hat{\tau}_{x}-\hat{k}_{y}\hat{\tau}_{y})s_{\bm{k}}$ or iv) chiral $d$-wave pairing $\hat{\Delta}_{\bm{k}}=-\Delta_{0}(\hat{k}_{x}^{2}-\hat{k}_{y}^{2})\hat{s}_{y}\hat{\tau}_{y}s_{\bm{k}}-2\Delta_{0}\hat{k}_{x}\hat{k}_{y}\hat{s}_{y}\hat{\tau}_{x}s_{\bm{k}}$. In both cases $s_{\bm{k}}=1$. Here, $\Delta_{0}$ is the gap amplitude in the presence of the AM field and can be suppressed [92, 99].
@@
where $\nu=\pm$ indicating two possible anisotropic energy shifts ($\pm h_{0}m_{\bm{k}}$) of subbands. From Eq. (2), it is easy to see that, for $\nu h_{0}m_{\bm{k}}=|\Delta_{0}s_{\bm{k}}|$, we get $E^{+}_{\bm{k}\nu}\neq 0$ and $E^{-}_{\bm{k}\nu}=0$, the latter signaling a zero energy bulk flat band akin to BFSs [23]. This is an example of a bulk subgap state induced by altermagnetism, and its detection can be carried out via conductance in the setup of Fig. 1(a). Notably, this setup also allows us to search for altermagnetism-induced subgap states at the surface, which can be directly accessed via conductance as we explain below.
To obtain the conductance, we employ the scattering approach, which is based on the wavefunctions of injected electrons with spin $\sigma=\uparrow,\downarrow$. Due to translational invariance along $y$, the wavefunctions are $\Psi_{\sigma}(x,k_{y})=\psi_{\sigma}(x)e^{ik_{y}y}$. In the normal side, $\psi_{\sigma}(x)$ is given by
$\psi_{\sigma}(x<0)=[\hat{A}_{\sigma}+a_{\sigma}\hat{B}_{\sigma}]e^{ik_{x}x}+b_{\sigma}\hat{A}_{\sigma}e^{-ik_{x}x}\,$, (3)
where $\hat{A}_{\uparrow}=\left(1,0,0,0\right)^{T}$, $\hat{A}_{\downarrow}=\left(0,1,0,0\right)^{T}$, $\hat{B}_{\downarrow}=\left(0,0,1,0\right)^{T}$ and $\hat{B}_{\uparrow}=\left(0,0,0,1\right)^{T}$, $k_{x}=k_{F}\cos\theta$, and $k_{y}=k_{F}\sin\theta$. The wave functions in the SC can be written in a similar way. Then, the coefficients $a_{\sigma}$ and $b_{\sigma}$ are obtained from the boundary conditions $\psi(0^{+})=\psi(0^{-})$, $\partial_{x}\psi(0^{+})-\partial_{x}\psi(0^{-})=(2mU_{0}/\hbar^{2})\psi(0)$. The angle-resolved charge conductance is found as [100] $\bar{\sigma}_{\sigma}^{e}=1+|a_{\sigma}|^{2}-|b_{\sigma}|^{2}$, which gives
$\bar{\sigma}_{\sigma}^{e}=\bar{\sigma}_{n}\frac{1+\bar{\sigma}_{n}|\Gamma_{\sigma}^{+}|^{2}+(\bar{\sigma}_{n}-1)|\Gamma_{\sigma}^{+}\Gamma_{\sigma}^{-}|^{2}}{|\Lambda-\eta(1-\bar{\sigma}_{n})\Gamma_{\sigma}^{+}\Gamma_{\sigma}^{-}|^{2}}\,$, (4)
@@
where $E_{\uparrow,\bm{k}_{\pm}}=E-\mathcal{M}_{\bm{k}_{\pm}}$ and $E_{\downarrow,\bm{k}_{\pm}}=E+\mathcal{M}_{\bm{k}_{\pm}}$, with $\bm{k}_{+}$, $\bm{k}_{-}$ are given by $\bm{k}_{\pm}=k_{F}\left(\pm\cos\theta,\sin\theta\right)$. Moreover, $\Lambda=1$ for $d_{x^{2}-y^{2}}$- and $d_{xy}$-wave SCs; $\Lambda=e^{2i\theta}$ for chiral $p$-wave SCs; $\Lambda=e^{4i\theta}$ for chiral $d$-wave SCs. Also, $\eta=1$ for spin-singlet SCs (cases i,ii,iv), while $\eta=-1$ for spin-triplet SCs (case iii). Hence, the charge conductance in the superconducting state is calculated as
$\sigma_{\mathrm{SC}}^{e}=\frac{e^{2}}{h}\int_{-\frac{\pi}{2}}^{\frac{\pi}{2}}\bar{\sigma}^{e}\cos\theta d\theta,$ (6)
where $\bar{\sigma}^{e}=\bar{\sigma}_{\uparrow}^{e}+\bar{\sigma}_{\downarrow}^{e}$ is the angle-resolved charge conductance. Later we normalize the conductance by $\sigma_{N}=(2e^{2}/h)\int\bar{\sigma}_{n}\cos\theta d\theta$.
@@
Bulk and surface flat bands by $d$-wave AMs in $d$-wave SCs.— We begin by analyzing the junction formed by $d_{x^{2}-y^{2}}$-wave SC with $d$-wave altermagnetism. Before delving into the conductance, we remind our discussion below Eq. (2), where we show that at $\nu h_{0}m_{\bm{k}}=|\Delta_{0}s_{\bm{k}}|$, one of the bands reaches zero energy, namely, $E_{\bm{k}\nu}^{-}=0$, while $E_{\bm{k}\nu}^{+}\neq 0$. Thus, when the AM and SC have the same symmetry ($m_{\bm{k}}=s_{\bm{k}}$) a zero-energy flat band emerges as a bulk subgap state entirely induced by altermagnetism and akin to BFSs. Indeed, this flat band is only possible due to the accidental disappearance of a gaped spectrum at $E=0$ while the superconducting gap still exists at $|E|<2\Delta_{0}$, maintaining superconducting coherence. As a result, the bulk density of states at zero energy becomes singular, similar to the divergent density of states at $E=\Delta_{0}$ for a conventional $s$-wave SC. This effect, and hence the flat band, can be detected via conductance, as we demonstrate in Fig. 2, where we plot the charge conductance showing nontrivial subgap states with $d$-wave AMs and $d$-wave SCs. The first feature to notice is that, for a $d_{x^{2}-y^{2}}$-wave SC with $d_{x^{2}-y^{2}}$-wave AM, the angle-resolved charge conductance $\bar{\sigma}^{e}(eV)$ at $h_{0}=\Delta_{0}$ is greatly enhanced at $eV=0$ as a result of the BFS ($E_{\bm{k}\nu}^{-}=0$) discussed above [Fig. 2(a)]. This enhancement gives rise to a zero-bias conductance peak (ZBCP) in the total charge conductance $\sigma_{\mathrm{SC}}^{e}(eV)$ at $h_{0}=\Delta_{0}$ [Fig. 2(b)]. The ZBCP stems from bulk BFSs and thus its width is very narrow [101]. By moving away from $eV=0$ and for $h_{0}\neq\Delta_{0}$, $\sigma_{\mathrm{SC}}^{e}(eV)$ develops a V-shape profile with spikes outside the gap, an effect that is similar to ferromagnetic SCs [39]. It is worth mentioning that when $d_{xy}$-wave AM and $d_{x^{2}-y^{2}}$-wave SC coexist, no zero energy bulk flat band can be found according to Eq. (2).
Figure 2: Conductance for a junction with various combinations of $d$-wave SCs and $d$-wave AMs. The AM is $d_{x^{2}-y^{2}}$-wave in panels (a-c) and $d_{xy}$-wave in (d). SC pairing is $d_{x^{2}-y^{2}}$-wave in (a), (b) and $d_{xy}$-wave in (c), (d). Panel (a) show angle-resolved conductance at $h_{0}=\Delta_{0}$. For all panels $Z=2$.
@@
Now we turn to $d_{xy}$-wave SCs. It is well-known that $d_{xy}$-wave pair potentials lead to zero-energy surface flat states without AMs [12]. Here we are interested in analyzing how AM can control the surface states and how to induce them in regimes not previously present. This is demonstrated in Figs. 2(c,d), where we plot the total charge conductance. For a $d_{x^{2}-y^{2}}$-wave AM and $d_{xy}$-wave SC, the total charge conductance $\sigma_{\mathrm{SC}}^{e}$ gets suppressed as the AM field increases, going from a sharp ZBCP at $h_{0}=0$ into an almost constant subgap conductance for $h_{0}>0$ whose size determines the strength of altermagnetism [Fig. 2(c)]. For a $d_{xy}$-wave AM and $d_{xy}$-wave SC, where the superconducting gap nodes are intact, the surface flat bands inherent to $d_{xy}$-wave SCs remain robust against the AM field when $h_{0}<\Delta_{0}$, see Fig. 2(d). This leads to a ZBCP in $\sigma_{\mathrm{SC}}^{e}(eV)$ for $h_{0}<\Delta_{0}$, but it reduces when $h_{0}>\Delta_{0}$ into an almost constant value within $|eV|<h_{0}-\Delta_{0}$. Further insights are obtained by looking at the zero-bias angle-resolved charge conductance for the $d_{xy}$-wave AM,
@@
Figure 3: Conductance for junctions with chiral $p$-wave SCs. AM is $d_{x^{2}-y^{2}}$ in (a)(c) and $d_{xy}$ in (b)(d). (a)(b) angle-resolved charge conductance at $h_{0}=\Delta_{0}$. (c)(d) total conductance for distinct values of $h_{0}$. $Z=2$ for all panels.
Surface curved bands by $d$-wave altermagnetism in chiral $p$-wave SCs.— Chiral $p$-wave pairing was once believed to describe superconductivity in Sr2RuO4 with chiral edge states [102, 103, 104], and a recent study even supports that Sr2RuO4 supports intrinsic AM order [105]. Motivated by these ideas, here we consider a one-band model with chiral $p$-wave pairing and $d$-wave altermagnetism; in the End Matter, we also address $p$-wave pairing with $g$-wave altermagnetism. For a $d_{x^{2}-y^{2}}$-wave AM, we calculate the chiral SABSs by imposing $\Gamma_{\sigma}^{+}\Gamma_{\sigma}^{-}=-e^{2i\theta}$ in Eq. (7),
$E_{\text{ABS}\nu}=\Delta_{0}\sin\theta+\nu h_{0}\cos 2\theta,$ (9)
with $\nu=\pm$. This reveals that the altermagnet field $h_{0}$ induces a spin splitting in the chiral SABSs $E_{\text{ABS}}^{\pm}$, which appear with a curved dispersion with $\theta$ and also revealed by angle-resolved conductance in Fig. 3(a). The splitting arises due $\Gamma_{\uparrow}^{+}\Gamma_{\uparrow}^{-}\neq\Gamma_{\downarrow}^{+}\Gamma_{\downarrow}^{-}$ in a $d_{x^{2}-y^{2}}$-wave AM and there are thus two species of SABSs as shown in Fig. 3(a). At certain angles found from $\Delta_{0}\sin\theta=\pm h_{0}\cos 2\theta$, however, the SABSs remain degenerate at zero energy [Fig. 3(a)], which affects their dispersion and makes them to acquire a quasi-flat band profile for a range of $\theta$. These quasi-flat bands then induce spikes in the total charge conductance whose separation for $h_{0}>\Delta_{0}/(2\sqrt{2})$ [Fig. 3(c)] is given by $\Delta_{0}^{2}/\left(4h_{0}\right)+2h_{0}$ from $\partial E_{\text{ABS}\nu}/\partial\theta=0$ and hence determines the strength of altermagnetism.